Solve for x
-2(x+7)+85=7x-9(x+3)

Answer by psbhowmick(465)

-2(x + 7) + 85 = 7x - 9(x + 3)
-2x - 14 + 85 = 7x - 9x - 27
-2x + 71 = -2x - 27
-2x + 2x = -71 - 27
0 = -98

So the equation is inconsistent.